Let G(x) = x5 + x4 + x2 + 1 be the polynomial generator(pattern), T = 101000110101110 be the message received by the destination node. Would the receiver accept the message? Why?


1
Expert's answer
2021-09-27T15:48:04-0400

Given polynomial is

"G(x) = x^5 +x^4+x^2+1"

Message from the destination node"(T)=101000110101110"

The Divisor polynomial bit "=1\times x^5 +1\times x^4+0\times x^3+1\times x^2+0\times x+1\times1"

"=110101"

Bits to be appended to the message = (Divisor polynomial bits -1)"=6-1= 5"

Now, append 5 zero to the message bit, hence"=10100011010111000000"

"10100011010111000000\/110101" Remainder =0000

Hence, CRC = 0000

Hence, we will not append any digit in the original message

So, the receiver will receive 101000110101110

Because there is not any error in the code, code is correct.